OpenStudy (anonymous):

converge or diverge sum_{2}^{infty} 1/(k ln k) use integral test

OpenStudy (experimentx):

\[ \sum_{k=2}^{\infty} \frac{1}{(k ln k)} \] Now that looks better.

OpenStudy (anonymous):

yes thanks

OpenStudy (turingtest):

\[\sum_{k=2}^{\infty}{1\over k\ln k}\]now it looks even better :! ok let's do it! (looking through notes)

OpenStudy (experimentx):

\[ \int_{2}^{\infty}\frac{1}{k \ln k} = ? \]

OpenStudy (anonymous):

u-sub ??

OpenStudy (anonymous):

idk

OpenStudy (turingtest):

yes, u-sub should work

OpenStudy (experimentx):

find out that, if that value is less than infinity then it should converge .... says wikipedia http://en.wikipedia.org/wiki/Convergence_test

OpenStudy (turingtest):

if the integral converges so does the series if the integral diverges, so does the series that's my understanding

OpenStudy (turingtest):

same thing as experimentX said I guess...

OpenStudy (turingtest):

@experimentX I'm gonna call you exX from now on, I'm lazy :)

OpenStudy (experimentx):

sure .. whatever

OpenStudy (anonymous):

i set my u = x and i got stuck

OpenStudy (experimentx):

no let u = ln x

OpenStudy (turingtest):

well that ain't the right sub actually u=x is never the right sub

OpenStudy (turingtest):

u=x is redundant you are just giving x a different name lol

OpenStudy (experimentx):

hahaha ... that s funny

OpenStudy (anonymous):

okay thats true

OpenStudy (experimentx):

well give it a try out ... alert us when you get stuck

OpenStudy (anonymous):

so i got \[\lim_{b \rightarrow \infty}\int\limits_{2}^{b} \ln(lnx)\]

OpenStudy (anonymous):

am i on the right track???

OpenStudy (turingtest):

yup, so does it diverge or converge?

OpenStudy (turingtest):

wait a minute... I see something weird...

OpenStudy (anonymous):

diverge

OpenStudy (anonymous):

because it goes to infinity

OpenStudy (turingtest):

do you see the same problem I do exX ? ln(2)<0 so ln(ln(2)) is undefined

OpenStudy (turingtest):

yeah, the other part definitely diverges, but what about evaluation at 2 ?

OpenStudy (anonymous):

ln(ln(2)) is just a constant i thought it shouldn't matter if i subtracted it from infinity

OpenStudy (experimentx):

let's ask the wolf first http://www.wolframalpha.com/input/?i=integrate+from+2+to+infinity+1%2F%28x+lnx%29 wolf says does not converge.

OpenStudy (experimentx):

I think it should be rather : ln(ln(inf)) - ln(ln(2)) ln 2 > 1 so lnln2 >1

OpenStudy (turingtest):

oh I'm trippin' ^^^

OpenStudy (turingtest):

yeah, so divergent

OpenStudy (anonymous):

cool

OpenStudy (turingtest):

ln2<1 not ln2<0 like I said befor

OpenStudy (experimentx):

Oo ... looks like i didn't see that.

OpenStudy (experimentx):

http://www.wolframalpha.com/input/?i=ln%28ln+%282%29%29 it's negative.

OpenStudy (turingtest):

but still defined and constant, so it's all good :)

OpenStudy (experimentx):

haha ... I forgot 2 < e